# Cognitive Foundations for Reasoning and Their Manifestation in LLMs

### v1

@@ 2.1 Reasoning Invariants: Computational Constraints
At Marr’s computational level, we investigate the fundamental constraints that an ideal solution to any reasoning problem must satisfy. Fodor’s (fodor1975language) language of thought hypothesis identifies these constraints: valid reasoning manipulates structured representations according to compositional rules, producing thoughts that are logically coherent (free from contradiction), compositional (complex meanings built from simpler parts), and productive (capable of generating unbounded novel inferences from finite primitives). We add conceptual processing from research on cognitive processing capacity and conceptual structures (halford1989cognitive; kholodnaya2016conceptual), which requires that reasoning operates over abstract relational structures rather than surface forms. These four invariants specify the computational constraints that define valid reasoning solutions.
@@
maintains consistency across reasoning steps and contexts (rips1983cognitive; thagard2002coherence). In the LEGO example, if the child simultaneously believes "this wing design is stable" and "this wing design will collapse," reasoning breaks down. The contradiction creates cognitive dissonance that must be resolved by revising the stability judgment, reinforcing the structure, or reconsidering the design (festinger1962cognitive; harmon2019introduction). This pressure to restore consistency shapes how reasoners revise mental models and update inferences. Humans cannot comfortably hold contradictory beliefs simultaneously; the drive to resolve inconsistencies is a defining feature of reasoning.
@@
enables building complex ideas from simpler components through rule-governed combination (fodor1975language; fodor2001language; fodor1988connectionism). The LEGO spaceship emerges from systematically combining individual bricks. The structure of the whole derives predictably from the properties of its parts and how they connect. Understanding "red cockpit with transparent dome" requires combining concepts of color, component type, and material in ways that preserve each element’s meaning. The meaning of a complex expression derives from the meanings of its constituents and their mode of combination (russin2024frege). This property drastically increases the learning efficiency and expressive power of the cognitive system, allowing finite mental resources to produce unbounded conceptual richness (frankland2020concepts; piantadosi2011learning).
@@
These four invariants are interconnected requirements rather than independent properties. Coherence constrains which compositional structures are valid. Productivity depends on compositional structure to generate novel thoughts systematically. Conceptual processing ensures reasoning operates over meaningful relationships rather than surface patterns. Together, these invariants define the computational constraints that valid reasoning must satisfy. However, invariants specify what reasoning must achieve without addressing the mechanisms that recognize when constraints are violated, select appropriate strategies, or monitor progress toward goals. These regulatory mechanisms constitute meta-cognitive controls.
@@ 2.2 Meta-Cognitive Controls: Executive Regulation
stands as the foundation: the capacity to assess one’s own knowledge state, capabilities, and the solvability of a task (neisser1988five; wicklund1979influence; leary2003evolution). Rochat (rochat2003five) describes this as "arguably the most fundamental issue in psychology, from both a developmental and evolutionary perspective". In the LEGO example, the child engages in internal assessment: "Am I good at building spaceships?" or "Do I know what a spaceship looks like?" This metacognitive evaluation enables strategic deployment of cognitive resources.
@@
These strategic choices are accompanied by heightgoal management: the process of setting, sequencing, and dynamically adjusting goals throughout reasoning (griffiths2019doing; dolan2013goals; cushman2015habitual). The abstract goal "build a spaceship" decomposes into manageable sub-goals: construct the main body, add wings, build the cockpit, and insert pilots. Goal management is inherently dynamic, corresponding to strategy shifts due to changes in context and self-awareness. While working on wings, the child might spot a clear brick perfect for a windshield, pausing the current task to insert a new sub-goal. Conversely, if attaching wheels proves too difficult, the child abandons this sub-goal to prioritize completing the rest of the ship.
@@ 2.4 Reasoning Operations: Transformation Procedures
Representation navigation includes operations that traverse knowledge and inference structures. heightForward chaining reasons from known facts toward goals: "These pieces connect this way; therefore this structure is possible" (huys2012bonsai). This operation is a key component of early production systems and expert system architectures, where the system begins with facts and applies rules iteratively until a conclusion is reached or a goal state is satisfied (newell1972human; davis1984origin) heightBackward chaining works from goals to prerequisites: "To make wings stable, I need reinforcement; what pieces provide that?" (park2017relative). This goal-driven approach is highly efficient in situations where the number of possible outcomes is large, but the goal is clearly defined (charniak1985introduction; newell1959report). heightBacktracking revisits earlier reasoning upon detecting errors (qin2025backtrack). This crucial cognitive control mechanism often relies on depth-first search, where, upon hitting a dead-end, the system returns to the most recent decision point to explore alternatives (Nilsson1971ProblemsolvingMI). In the LEGO example, the child uses forward chaining to build, backward chaining to plan, and backtracking when a design fails. These navigation operations determine the path through problem space.
These operations do not occur in isolation. In the LEGO example, representation selection chooses hierarchical decomposition, forward chaining builds the body, verification checks that pieces connect properly, backtracking returns to an earlier design choice when the structure wobbles, restructuring introduces new design ideas, pattern recognition suggests a previously worked design, and abstraction extracts a principle for future builds. The operations form an integrated toolkit, flexibly deployed under meta-cognitive control to manipulate representations while satisfying invariants. Together with the representational structures from Section 2.3, they constitute Marr’s algorithmic level. Table 2 presents our complete taxonomy: invariants define computational constraints, controls select and monitor processes, representations encode knowledge structures, and operations transform those structures. This framework provides the analytical vocabulary for examining how these 28 reasoning behaviors manifest in human versus LLM reasoning. Section 3 presents our investigation of behavioral presence, patterns, and their relationship to reasoning success across diverse problem types.
@@ 3.1.1 Data Collection
Total (Text Data) 10,612 Total (Text Models) 14 148,568
Audio BLAB (ahia2025blabbrutallylongaudio) 417 Qwen/Qwen3-Omni-30B-A3B-Thinking qwen3 1 4,917
MMAR (ma2025mmar) 888
@@
Image Zebra-CoT (li2025zebracot) 18,000 GPT 4.1 and Gemini 2.5 pro (Only used for refinement) 2 18,000
Total (Image Data) 18,000 Total (Image Models) 2 18,000
Grand Total (All Data) 33,529 Grand Total (All Models) 17 171,485
@@
We analyze 14 open-weight text reasoning models spanning multiple architecture families and training paradigms. These include: Qwen3 hybrid models with thinking mode, R1-Distilled models built on Qwen 2.5, R1-Distilled models built on Llama 3, several community-developed reasoning-tuned models (DeepScaleR-1.5B-Preview, s1.1-32B, OpenThinker-32B, DeepHermes-3-Llama-3-8B-Preview), and the frontier DeepSeek-R1 (671B), which also serves as the teacher model for R1-Distilled variants. Text-only reasoning problems consist of 10,612 questions sampled from the GeneralThought (RJT1990-GeneralThoughtArchive) and ClaimSpect (kargupta-etal-2025-beyond) datasets. The task label in the dataset is used to down-sample extremely common tasks (e.g., arithmetic, simple logic) to maintain a balanced representation. We use released DeepSeek-R1 traces from GeneralThought due to compute constraints. While GeneralThought provides a wide variety of tasks and domains, the dataset primarily focuses on verifiable tasks and consequently lacks problems of the type Dilemma [^2] Dilemma type: resolving situations with contradictory positions and no clear satisfactory solution (Appendix A.2).. Thus, we convert 290 real-world biomedical and geopolitical "nuanced" claims from ClaimSpect into questions to supplement the text-reasoning category.
@@
For image reasoning problems, we directly use the reasoning traces in Zebra-CoT due to its comprehensive task type coverage and curated reasoning traces. In Zebra-CoT, real-world traces are sourced from online math, physics, coding, and chess datasets. Synthetic traces are created by generating or sourcing images online and writing reasoning templates, then using frontier VLMs (Gemini-2.5 and GPT-4.1) to refine them into more diverse and coherent reasoning traces (li2025zebracot). Additionally, this allows us to analyze cognitive elements in synthetic training data beyond raw model reasoning outputs. We sample 1,000 question-reasoning pairs from each task type to obtain 18,000 curated reasoning traces. Across all three modalities, we collect 171,485 model reasoning traces for detailed analysis.
@@
We collected a small set of human reasoning traces, as qualitative reference points for comparison with LLM-generated reasoning. We recruited 18 human participants to solve a small subset of the GeneralThought dataset while recording their reasoning. These human traces overlap with a subset of the LLM evaluation set and are intended to illustrate how key behaviors manifest in natural human reasoning, rather than to establish a full human benchmark.
@@
Each reasoning trace was annotated separately by two different human annotators, giving each reasoning trace a score from a three-level scoring rubric (0=absent, 1=partially present, 2=present) across each behavior in our 28-behavior taxonomy. Scores from different annotators were aggregated via min-pooling, ensuring that estimates of cognitive elements are conservative. These annotation data were later used for iteratively refining the automatic span annotation prompts.
@@ 3.1.2 Fine-Grained Cognitive Element Annotation
Each reasoning trace is annotated for the presence of behaviors that embody the cognitive elements from the 28-behavior taxonomy introduced in Section 2 on the span-level. The annotation process identifies specific text segments that demonstrate each cognitive capability, enabling precise localization of behaviors within reasoning processes.
@@ 3.1.3 Problem Type Classification & Response Evaluation
We extend the problem-solving taxonomy in jonassen2000toward to classify cognitive tasks in our dataset. jonassen2000toward defines problem solving as a goal-directed cognitive activity that transforms an initial state into a desired goal state through systematic reasoning, proposing a typology of 11 problem types: logical problems, algorithms, story problems, rule-using, decision-making, troubleshooting, diagnosis-solution, strategic performance, case analysis, design, and dilemmas. jonassen2015all further characterizes ten out of the eleven problems types along a continuum from well-structured (clear goals, known solution paths, predictable outcomes) to ill-structured (ambiguous goals, multiple solution paths, uncertain outcomes). However, not all tasks in the GeneralThought dataset require goal-directed transformation. We add two categories to capture tasks outside this paradigm: factual recall (retrieving stored knowledge without reasoning) and creative/expressive tasks (generating novel content judged by originality or aesthetic quality rather than convergence to a predetermined solution). This yields a 13-category taxonomy (Appendix A.2) spanning the full spectrum of cognitive demands in our datasets.
@@
3.1.4 Cognitive Structure Extraction
Beyond assessing each behavior’s presence within a reasoning trace, our span-level annotation approach enables fine-grained, quantitative analysis of behaviors and their interdependencies. Specifically, we seek to analyze how a reasoner subconsciously structures and sequences specific behaviors throughout their reasoning process. In order to extract this structure, we choose to encode the reasoning trace $t$ as a heterogeneous transition graph $G$, where each node represents a single behavior from our 28-behavior taxonomy, and each edge can either represent a hierarchical (CONTAINS), sequential (NEXT), or parallel (PAR) behavior relationships. We compute a weight $w$ for each node and edge based on its normalized pointwise mutual information (NPMI) score. We calculate the NPMI using the individual and joint probabilities of (a) behavior $b$ occurring in a reasoning trace $t$ and (b) trace $t$ successfully resulting in a correct answer.
In order to construct $G$, we first sort all annotated behavior spans within trace $t$ by their start positions, with ties broken by span length in descending order. For each pair of behaviors $(b_{a},b_{b})$ with corresponding spans $s_{a},e_{a}]$ and $s_{b},e_{b}]$ where $s_{a}\leq s_{b}$, we determine their relationship type through a multi-stage classification process:
@@ 3.1.4 Cognitive Structure Extraction
    PAR: Whether the Manhattan distance between the spans ($|s_{b}-s_{a}|+|e_{b}-e_{a}|$) falls below a threshold $\tau_{par}$ ($b_{a}$ and $b_{b}$ occur nearly simultaneously).
@@
We set $\tau_{par}=20$ (characters) and $\tau_{overlap}=0.8$. To ensure graph sparsity and capture only direct sequential dependencies, we apply a refinement step: for each behavior $b_{a}$, we retain only the first NEXT edge to an immediately subsequent non-overlapping behavior, filtering out transitive connections to more distant behaviors. This process yields a directed graph that preserves both the hierarchical decomposition and temporal ordering of reasoning behaviors within each trace.

### v2

@@ 2.1 Reasoning Invariants: Properties & Goals
We first investigate the fundamental properties of reasoning, which are computational goals that constrain the ideal solution to any reasoning problems. Fodor’s (fodor1975language) language of thought hypothesis identifies these properties: valid reasoning manipulates structured representations according to compositional rules, producing thoughts that are logically coherent (free from contradiction), compositional (complex meanings built from simpler parts), and productive (capable of generating unbounded novel inferences from finite primitives). We add conceptual processing from research on cognitive processing capacity and conceptual structures (halford1989cognitive; kholodnaya2016conceptual), which requires that reasoning operates over abstract relational structures rather than surface forms. These four invariants specify the computational goals that define and constrain valid reasoning solutions.
@@
maintains consistency across reasoning steps and contexts (rips1983cognitive; thagard2002coherence). In the LEGO example, if the child simultaneously believes “this wing design is stable” and “this wing design will collapse,” reasoning breaks down. The contradiction creates cognitive dissonance that must be resolved by revising the stability judgment, reinforcing the structure, or reconsidering the design (festinger1962cognitive; harmon2019introduction). This pressure to restore consistency shapes how reasoners revise mental models and update inferences. Humans cannot comfortably hold contradictory beliefs simultaneously; the drive to resolve inconsistencies is a defining feature of reasoning.
@@
enables building complex ideas from simpler components through rule-governed combination (fodor1975language; fodor2001language; fodor1988connectionism). The LEGO spaceship emerges from systematically combining individual bricks. The structure of the whole derives predictably from the properties of its parts and how they connect. Understanding “red cockpit with transparent dome” requires combining concepts of color, component type, and material in ways that preserve each element’s meaning. The meaning of a complex expression derives from the meanings of its constituents and their mode of combination (russin2024frege). This property drastically increases the learning efficiency and expressive power of the cognitive system, allowing finite mental resources to produce unbounded conceptual richness (frankland2020concepts; piantadosi2011learning).
@@
These four invariants are interconnected requirements rather than independent properties. Coherence constrains which compositional structures are valid. Productivity depends on compositional structure to generate novel thoughts systematically. Conceptual processing ensures reasoning operates over meaningful relationships rather than surface patterns. Together, these invariants define the computational goals and constraints that valid reasoning must satisfy. However, invariants specify what reasoning must achieve without addressing the mechanisms that recognize when constraints are violated, select appropriate strategies, or monitor progress toward goals. These regulatory mechanisms constitute meta-cognitive controls.
@@ 2.2 Meta-Cognitive Controls: Executive Regulation
stands as the foundation: the capacity to assess one’s own knowledge state, capabilities, and the solvability of a task (neisser1988five; wicklund1979influence; leary2003evolution). Rochat (rochat2003five) describes this as “arguably the most fundamental issue in psychology, from both a developmental and evolutionary perspective.” In the LEGO example, the child engages in internal assessment: “Am I good at building spaceships?” or “Do I know what a spaceship looks like?” This metacognitive evaluation enables strategic deployment of cognitive resources.
@@
These strategic choices are accompanied by heightgoal management: the process of setting, sequencing, and dynamically adjusting goals throughout reasoning (griffiths2019doing; dolan2013goals; cushman2015habitual). The abstract goal “build a spaceship” decomposes into manageable sub-goals: construct the main body, add wings, build the cockpit, and insert pilots. Goal management is inherently dynamic, corresponding to strategy shifts due to changes in context and self-awareness. While working on wings, the child might spot a clear brick perfect for a windshield, pausing the current task to insert a new sub-goal. Conversely, if attaching wheels proves too difficult, the child abandons this sub-goal to prioritize completing the rest of the ship.
@@ 2.4 Reasoning Operations: Transformation Procedures
Representation navigation includes operations that traverse knowledge and inference structures. heightForward chaining reasons from known facts toward goals: “These pieces connect this way; therefore this structure is possible” (huys2012bonsai). This operation is a key component of early production systems and expert system architectures, where the system begins with facts and applies rules iteratively until a conclusion is reached or a goal state is satisfied (newell1972human; davis1984origin) heightBackward chaining works from goals to prerequisites: “To make wings stable, I need reinforcement; what pieces provide that?” (park2017relative). This goal-driven approach is highly efficient in situations where the number of possible outcomes is large, but the goal is clearly defined (charniak1985introduction; newell1959report). heightBacktracking revisits earlier reasoning upon detecting errors (qin2025backtrack). This crucial cognitive control mechanism often relies on depth-first search, where, upon hitting a dead-end, the system returns to the most recent decision point to explore alternatives (Nilsson1971ProblemsolvingMI). In the LEGO example, the child uses forward chaining to build, backward chaining to plan, and backtracking when a design fails. These navigation operations determine the path through problem space.
These operations occur on the reasoning representations. In the LEGO example, representation selection chooses hierarchical decomposition, forward chaining builds the body, verification checks that pieces connect properly, backtracking returns to an earlier design choice when the structure wobbles, restructuring introduces new design ideas, pattern recognition suggests a previously worked design, and abstraction extracts a principle for future builds. The operations form an integrated toolkit, flexibly deployed under meta-cognitive control to manipulate representations while satisfying invariants. Together with the representational structures from Section 2.3, they constitute Marr’s algorithmic level. Table 2 presents our complete taxonomy: invariants define computational goals and constraints, controls select and monitor processes, representations encode knowledge structures, and operations transform those structures. This framework provides the analytical vocabulary for examining how these 28 cognitive elements manifest in human versus LLM reasoning behaviors. Section 3 presents our investigation of behavioral presence, patterns, and their relationship to reasoning success across problem types.
@@ 3.1.1 Data Collection
Olmo 3 (7B, 32B) (olmo3) 2 21,224
Total (Text Data) 10,612 Total (Text Models) 16 169,792
Audio BLAB (ahia2025blabbrutallylongaudio) 417 Qwen3-Omni-30B-A3B-Thinking 1 4,917
MMAR (ma2025mmar) 888 qwen3
@@
Image Zebra-CoT (li2025zebracot) 18,000 GPT 4.1 and Gemini 2.5 pro (Only used for refinement, collapse to 1 model) 1 18,000
Total (Image Data) 18,000 Total (Image Models) 1 18,000
Grand Total (All Data) 33,529 Grand Total (All Models) 18 192,709
@@
We analyze 16 open-weight text reasoning models spanning multiple architecture families and training paradigms. These include: Qwen3 hybrid models with thinking mode, R1-Distill models built on Qwen 2.5, R1-Distill models built on Llama 3, Olmo 3 models with thinking mode, several community-developed reasoning-tuned models (DeepScaleR-1.5B, s1.1-32B, OpenThinker-32B, DeepHermes-3-Llama-3-8B-Preview), and the frontier DeepSeek-R1 (671B), which also serves as the teacher model for R1-Distilled variants. Text-only reasoning problems consist of 10,612 questions sampled from the GeneralThought (RJT1990-GeneralThoughtArchive) and ClaimSpect (kargupta-etal-2025-beyond) datasets. The task label in the dataset is used to down-sample extremely common tasks (e.g., arithmetic, simple logic) to maintain a balanced representation. We use released DeepSeek-R1 traces from GeneralThought due to compute constraints. While GeneralThought provides a wide variety of tasks and domains, the dataset primarily focuses on verifiable tasks and consequently lacks problems of the type Dilemma [^2] Dilemma type: resolving situations with contradictory positions and no clear satisfactory solution (Appendix A.2).. Thus, we convert 290 real-world biomedical and geopolitical “nuanced” claims from ClaimSpect into questions to supplement the text-reasoning category.
@@
For image reasoning problems, we directly use the reasoning traces in Zebra-CoT due to its comprehensive task type coverage and curated reasoning traces. In Zebra-CoT, real-world traces are sourced from online math, physics, coding, and chess datasets. Synthetic traces are created by generating or sourcing images online and writing reasoning templates, then using frontier VLMs (Gemini-2.5 and GPT-4.1) to refine them into more diverse and coherent reasoning traces (li2025zebracot). Additionally, this allows us to analyze cognitive elements in synthetic training data beyond raw model reasoning outputs. We sample 1,000 question-reasoning pairs from each task type to obtain 18,000 curated reasoning traces. Across all three modalities, we collect 192,709 model reasoning traces for detailed analysis.
@@
We collected a small set of human reasoning traces, as qualitative reference points for comparison with LLM-generated reasoning. We recruited 18 human participants to solve a small subset of the GeneralThought dataset while recording their reasoning. These human traces overlap with a subset of the LLM evaluation set and are intended to illustrate how key elements manifest in natural human reasoning, rather than to establish a full human benchmark.
@@
Each reasoning trace was annotated separately by two different human annotators, giving each reasoning trace a score from a three-level scoring rubric (0=absent, 1=partially present, 2=present) across each element in our 28-element taxonomy. Scores from different annotators were aggregated via min-pooling, ensuring that estimates of cognitive elements are conservative. These annotation data were later used for iteratively refining the automatic span annotation prompts.
@@ 3.1.2 Fine-Grained Cognitive Element Annotation
Each reasoning trace is annotated for the presence of behaviors that embody the cognitive elements from the 28-element taxonomy introduced in Section 2 on the span-level. The annotation process identifies specific text segments that demonstrate each cognitive capability, enabling precise localization of behaviors within reasoning processes.
@@ 3.1.3 Problem Type Classification & Response Evaluation
We extend the problem-solving taxonomy in jonassen2000toward to classify cognitive tasks in our dataset. jonassen2000toward defines problem-solving as a goal-directed cognitive activity that transforms an initial state into a desired goal state through systematic reasoning, proposing a typology of 11 problem types: logical problems, algorithms, story problems, rule-using, decision-making, troubleshooting, diagnosis-solution, strategic performance, case analysis, design, and dilemmas. jonassen2015all further characterizes ten out of the eleven problem types along a continuum from well-structured (clear goals, known solution paths, predictable outcomes) to ill-structured (ambiguous goals, multiple solution paths, uncertain outcomes). However, not all tasks in the GeneralThought dataset require goal-directed transformation. We add two categories to capture tasks outside this paradigm: factual recall (retrieving stored knowledge without reasoning) and creative/expressive tasks (generating novel content judged by originality or aesthetic quality rather than convergence to a predetermined solution). This yields a 13-category typology (Appendix A.2) spanning the full spectrum of cognitive demands in our datasets.
@@
3.1.4 Reasoning Structure Construction
Beyond assessing each cognitive element’s presence within a reasoning trace, our span-level annotation approach enables fine-grained, quantitative analysis of their behavioral manifestations and interdependencies. Specifically, we seek to analyze how a reasoner subconsciously structures and sequences specific elements throughout their reasoning process. To construct this structure, we choose to encode the reasoning trace $t$ as a heterogeneous transition graph $G$, where each node represents a single element from our 28-element taxonomy, and each edge can represent a hierarchical (CONTAINS), sequential (NEXT), or parallel (PARALLEL) relationship. We compute a weight $w$ for each node and edge based on its normalized pointwise mutual information (NPMI) score. We calculate the NPMI using the individual and joint probabilities of (a) element $b$ manifesting in a reasoning trace $t$ and (b) trace $t$ successfully resulting in a correct answer.
To construct $G$, we first sort all annotated spans within trace $t$ by their start positions, with ties broken by span length in descending order. For each pair of elements $(b_{a},b_{b})$ with corresponding spans $s_{a},e_{a}]$ and $s_{b},e_{b}]$ where $s_{a}\leq s_{b}$, we determine their relationship type through a multi-stage classification process:
@@ 3.1.4 Reasoning Structure Construction
    PARALLEL: Whether the Manhattan distance between the spans ($|s_{b}-s_{a}|+|e_{b}-e_{a}|$) falls below a threshold $\tau_{par}$ ($b_{a}$ and $b_{b}$ occur nearly simultaneously).
@@
We set $\tau_{par}=20$ (characters) and $\tau_{overlap}=0.8$. To ensure graph sparsity and capture only direct sequential dependencies, we apply a refinement step: for each element $b_{a}$, we retain only the first NEXT edge to an immediately subsequent non-overlapping element, filtering out transitive connections to more distant elements. This process yields a directed graph that preserves both the hierarchical decomposition and temporal ordering of manifestations of cognitive elements within each trace. Each node and edge in graph $G$ is weighted by its NPMI score, quantifying the strength of association between that behavioral element and trace success.
Prototypical reasoning structure extraction. For each problem type, we extract $\hat{G}$ representing the most frequently deployed behavioral patterns across all traces. We aggregate transition statistics across all reasoning graphs $\{G_{t}\}$ for that problem type, computing edge occurrence frequencies $f(b_{\text{curr}},b_{\text{next}},e_{\text{type}})=|\{\text{traces containing this edge}\}|/|\{\text{total traces}\}|$. Starting from the most frequent initial element $b_{\text{start}}=\operatorname*{arg\,max}_{b}P(b\text{ appears first})$, we construct $\hat{G}$ through greedy forward search: at each step from $b_{\text{curr}}$, we select the outgoing edge $(b_{\text{curr}},b_{\text{next}},e_{\text{type}})$ with maximum occurrence frequency among unvisited targets. We maintain a visited set to ensure acyclicity and continue until reaching $|V_{\max}|$ nodes or exhausting valid edges. This yields a prototype structure capturing how models typically organize the cognitive elements.
Successful reasoning structure extraction. To identify behavioral patterns characteristic of successful traces, we extract $G^{*}$ by selecting high-NPMI transitions from correct solutions. We begin from $b_{\text{start}}=\operatorname*{arg\,max}_{b}\text{NPMI}(b,\text{success})$ among elements that initiate successful traces. At each step from $b_{\text{curr}}$, we select the outgoing edge $(b_{\text{curr}},b_{\text{next}},e_{\text{type}})$ with maximum NPMI score among unvisited targets, maintaining acyclicity through a visited set. The process terminates when: (1) reaching $|V_{\max}|$ nodes, (2) no valid outgoing edges exist, or (3) all remaining edges have non-positive NPMI scores. This produces $G^{*}$ representing behavioral patterns strongly associated with correct reasoning for that problem type.